# Benchmarking Quantum Computers: Towards a Standard Performance Evaluation Approach

### v3

@@ III.3.3 Benchmarks in NISQ Era
QPack: The QPack benchmark utilizes Quantum Approximate Optimization Algorithms (QAOA) to evaluate the effectiveness of quantum computers in solving problems such as MaxCut, dominating set, and the traveling salesman problem QPACK2021 . This benchmark assesses various aspects of a quantum computer, including runtime, output accuracy, and scalability. The performance of the quantum computer is assessed by contrasting the solution obtained through QAOA execution on the quantum computer with that obtained using a classical exact algorithm. Consequently, the size of the problem that can be handle is constrained by the computational capacity of the classical computer. The QPack benchmark has been used to compare different quantum computing platforms, such as superconducting and trapped ion devices QPACK_ .
@@
Quantum LINPACK: Inspired by the classical LINPACK benchmark, the Quantum LINPACK benchmark assesses the performance of a quantum computer based on its ability to solve linear equations system using pseudo-dense random matrices. QuantumLINPACK . The creators of this benchmark introduce an input model known as the RAndom Circuit Block-Encoded Matrix (RACBEM), which serves as an appropriate generalization of a dense random matrix in the context of quantum computing. The success of the benchmark is measured by comparing the output vector to the ideal solution, incurring an exponential classical cost.
@@
QASMBench: The QASMBench is a benchmark that includes a set of quantum circuits and metrics for characterizing circuit properties both before and after execution on a quantum processor li2022qasmbench . The set of quantum circuits are frequently used in various fields, such as chemistry, simulation, linear algebra, search algorithms, optimization, arithmetic, machine learning, fault tolerance, cryptography, and more QuantumAlgorithms . QASMBench is divided into three categories: small-scale, medium-scale, and large-scale, each defined by the number of qubits used. As an illustration, in the small-scale scenario, the performance of quantum processes is assessed by executing quantum algorithms using 1-5 qubits such as Quantum Walk, Linear Equation, Search, and optimization, among others. In the medium-scale and large-scale scenarios, the performance of quantum processes is evaluated by running quantum algorithms using 6-15 and more than 15 qubits, respectively. The authors propose a set of circuit evaluation metrics: Circuit Width, Circuit Depth, Gate Density, Retention lifespan, Measurement Density, and Entanglement Variance, which are indicators that allow us to quantify the performance of a certain quantum circuit executed on a quantum processor. QASMBench has been assessed on IBM-Q machines, and the effectiveness of the quantum circuits is determined by computing the Hellinger distance.
@@
SupermarQ Suite: SupermarQ Suite is a benchmark that includes a set of applications that assess the performance of a quantum computer based on its ability to generate entanglement between qubits through the GHZ algorithm, demonstrate the quantum nature of the system through the Mermin-Bell test, solve combinatorial optimization problems through the quantum approximate optimization algorithm, find the ground state energy of the one-Dimensional Transverse Field Ising Model (1D TFIM) using the Variational Quantum Eigensolver algorithm (VQE) and simulate the time evolution of the 1D TFIM Hamiltonian using a quantum circuit via Trotterization for a specific number of time steps 9773202 . The performance of each application is assessed through different scoring function, comparing experimental outcomes with predefined ideal results. As these ideal results are known in advance, the scalability of this benchmark is ensured. The SupermarQ Suite has been employed to assess the performance of both superconducting and trapped ion devices.
@@
QED-C Benchmark: The Quantum Economic Development Consortium (QED-C) introduced the QED-C benchmark to assess the efficiency of quantum computers lubinski2023applicationoriented . This benchmark comprises various algorithms, including Deutsch-Josza, Grover’s Search, VQE, Phase Estimation, Bernstein-Vazirani, Quantum Fourier Transform, and other algorithms. Unlike the Quantum Volume metric, which only uses square circuits, the QED-C benchmark incorporates circuits that vary in both depth and width. It includes circuits with high depth and low width, as well as circuits with low depth and high width. The QED-C benchmark evaluates computer performance by assessing fidelity, which involves comparing the ideal probability distribution with the experimental probability distribution. Therefore, this approach is limited by its requirement for exponentially costly classical computations to evaluate the performance of quantum computers.
@@
Chemistry simulation: The Chemistry Simulation benchmark is designed to assess the performance of quantum computers by the electronic structure calculation of molecules using variational methods QuantumChemistry . The variational quantum eigensolver (VQE) algorithm has been used to calculate electronic structures using various quantum computing architectures both photonic VQE-Photonic , superconducting circuits PhysRevX.6.031007 and trapped ions PhysRevX.8.031022 . This benchmark assesses the performance of quantum computers by using ground state energy as a metric, comparing results obtained on classical computers with those derived using the VQE algorithm on quantum computers.
We have introduced the main proposals of metrics and benchmarks in the literature. Table 1 analyzes metrics, evaluating which quality attributes they fulfill according to the framework introduced in Sec. II. We emphasize that we evaluate the quality attributes fulfilled by metrics in the NISQ era; Therefore, these evaluations are not valid for the post-NISQ era. We hope that as quantum processors evolve, benchmarks will adapt to each era.
@@
Number of (Working Connected) Qubits Scalability
@@
(Average) Decoherence Times $\times$ - $\times$ Quality
@@
Quantum Volume $\times$ $\times$ $\times$ $\times$ Scalability / Quality
Q-Score $\times$ $\times$ $\times$ Scalability / Quality
CLOPS $\times$ $\times$ $\times$ $\times$ Speed
@@
Cross Entropy Benchmarking (XEB) $\times$ - Quality
@@
Table 1: Summary of quantum computing metrics in NISQ era. In this table, we show some of the main metrics that have been proposed to measure properties related to the performance of a quantum computer, as presented in subsection III.3.2. We assess whether these metrics satisfy the quality attribute proposed in the literature for an ideal performance metric, as discussed in subsection II.3, although some of them are borderline. The check mark indicates that the metric satisfies the attributes; otherwise, we use the cross mark. The dash mark indicates that it is complicated to define if the metric satisfies the criterion because we cannot characterize it formally. Performance definitions: Scalability refers to the size of the processor; Quality evaluates how accurately the quantum processor performs the tasks it was designed or programmed to execute; Speed concerns the time required to process a certain task.
IV Roadmap for Quantum Benchmarking
@@
IV.1 Guidelines for the benchmark practitioner
Here, we present a set of high-level guidelines to assist benchmark designers in making informed decisions when creating benchmarks that accurately reflect the performance of the assessed quantum devices. These guidelines build upon existing industry standards and take into account insights from classical benchmarking sources (as elaborated in Sec. II), being system-independent:
@@ IV.1 Guidelines for the benchmark practitioner
Guideline 2. Benchmarks must uphold the key quality attributes: relevance, reproducibility, fairness, verifiability, and usability. In Subsec. II.1, we introduced a set of quality attributes initially suggested as guidelines for classical benchmark designers. We aim to extend the application of these attributes to the field of quantum computing devices. Relevance aids in defining the specific area within the total behavior space targeted by our benchmark proposal and clarifies how accurately it captures this domain. Fairness is key for ensuring impartial comparisons across diverse gate-sets and processor architectures. Reproducibility and usability guarantee that results obtained by a particular device running the benchmark can be verified by other vendors, fostering trust among stakeholders. Finally, verifiability is instrumental in confirming the correctness of the results. A benchmark should be self-verifiable, to enable other parties to repeat the executions and obtain similar results. One of our key proposals and recommendations is that any newly defined benchmark should be accompanied by an analysis that clearly identifies which properties it satisfies and which it does not, along with a justification for the relevance or irrelevance of each property in the given context. This practice will help the community understand the appropriate contexts in which these benchmarks or metrics should be applied, as well as highlight key limitations that future developments should aim to address. As an illustrative example, we briefly analyze one of the benchmarks previously discussed in Subsec. III.3.3: the QED-C Benchmark. This benchmark, or benchmark suite, has been continuously updated to incorporate a variety of application-oriented benchmarks for quantum computers. The final output of the benchmark can be interpreted either as a vector comprising the results of multiple specified metrics, or as a collection of individual benchmarks, each with its own set of metrics. Regardless of the interpretation, the resulting analysis remains largely consistent. In this analysis, we treat it as a benchmark suite. Some components of the suite, such as those based on Shor or Grover algorithms, are not relevant in the current technological era, a point acknowledged in the original guidelines. Others, such as benchmarks based on the Variational Quantum Eigensolver (VQE), remain pertinent. In principle, the benchmarks are reproducible, provided that the corresponding metrics are reported alongside detailed information about all parameters and device calibration. They are also fair, as they do not introduce biases tied to specific hardware architectures or artificial constraints. The methodology is well-defined and established, contributing to the overall verifiability of the suite. However, some benchmarks, such as those involving Hamiltonian simulation, do not scale efficiently to larger qubit counts due to their reliance on classical simulation. This reliance renders them unverifiable for systems with many qubits, as classical resources become insufficient to reproduce or validate the results. Moreover, not all benchmarks in the suite are usable at scale; for example, in cases like Hamiltonian simulation, the required number of measurement shots increases significantly with the number of qubits, making them unsuitable for systems with many qubits.
@@
Guideline 3. Benchmarks must offer peak and base performance results. The classical baseline debate with arguments against and for an extended use of optimization flags can be extended to quantum devices. We encourage benchmarks practitioners to follow a similar approach as the final compromise established for SPEC in January 1994: ”Every SPEC CPU benchmark measurement has to measure performance with a restricted baseline set of flags, and optionally with a more extended peak set of flags”SPECHPGBenchmarksHistory . Having both metrics, peak and base, allows different vendors to share a common baseline for comparison that represents also the average non expert consumer, while enabling also the publication of the peak performance with optimised results leveraging expert knowledge over the architectures.
@@
Guideline 4. Benchmarks metrics must uphold the key quality attributes: practical, repeatable, reliable, and consistent. They should also preferably be independent and linear. As discussed in Sec. II, for a metric to be useful in build a good benchmark it must present attributes that correlate later with quality benchmark attributes. Additionally, while linearity and independence are desirable metric attributes, they are not strictly required for effectively capturing the performance of a quantum device. Regarding benchmarks, we encourage that each proposal of a new metric includes an analysis of the properties it satisfies as well as those it does not.
@@
IV.2 Benchmarking structure
@@
Device verification is a crucial process to ensure that the device is working properly. Just as vendors in classical computing certify the validity of their products before consumer release, similar steps must be taken in the quantum computing field. However, the techniques for verification differ due to the fundamentally different nature of quantum computers, rooted in quantum mechanics. Since we are in NISQ era, the concepts of verification and benchmarking overlap because we can not ensure that a quantum computer will work properly. An example of device verification method is the cross-entropy benchmarking (XEB), which has been employed to verify quantum processors based on superconducting circuits Neill_2018 ; Arute2019 .
@@
  • Device and routine execution verification. Post-NISQ era, a significant challenge we must address, which is becoming increasingly apparent with advancements in available quantum computers even within NISQ, is verifying the correctness of the results they produce. As mentioned earlier in Sect. III, we know that the problem to verify if every language in BQP admits an interactive proof where the prover is in BQP too, and the verifier is in BPP is still opened AaronsonPrize . This question has been thoroughly examined within the field, even extending to inquiries about the falsifiability of quantum mechanics as a theory aharonov2012quantum . Still, it is not the only approach explored for verifiability so far in the field, with other protocols explores with single-provers and multi-provers either quantum or classical Gheorghiu2018 .
@@
  • Definition of good metrics. In this study, we have introduced in Subsec. II.3 the quality attributes that good metrics should ideally meet. Additionally, we have examined various proposals put forth by researchers in the field in Subsec. III.3. However, we contend that further work is necessary to precisely define the relevance of the metrics. In other words, we need to determine which aspects of the devices behavior space are being captured by these measurements. Establishing a correlation between metrics and their actual relevance to performance is crucial for obtaining diverse perspectives of the device, and a thorough assessment. Throughout the NISQ era, prioritizing metrics focused on scalability, accuracy, and quality can better guide the development of robust and practically useful quantum devices. While gate speed remains a relevant and informative metric—especially given the variability in gate durations across different technological platforms—we consider it complementary to those assessing overall system performance and reliability. In our view, scalability, accuracy, and quality represent more immediate challenges for current quantum computing architectures. Therefore, speed-related metrics are valuable, but should be regarded as secondary to those that capture the system’s computational potential and robustness. However, speed may become a necessary performance indicator as we transition towards the Fault Tolerant era or implement additional partial quantum error correction methods.
@@
  • Standarise transpilation and compilation process. While we have discussed in detail the problematic regarding the transpilation, which can be mapped to the idea of how to optimise the intermediate representation (IR) language descriptor for quantum programs (e.g. OpenQASM cross2017open or QASM 1580386 ), we have not detailed the mapping into the physical layer. This aspect of the compilation and transpilation funnel relates to the Quantum Physical Operations Language (QPOL) 1580386 . Due to the lack of a standard architecture for the development of the quantum technologies and the complex nature of simplifying the intermediate representation (IR) of circuits, much of the research focus has revolved around this aspect. In the future, we might have to include the exploration of QPOL optimization techniques through standard compilers within the industry. Even today, transpilation alone is witnessing a surge of initiatives aiming to delve deeper into its complexities. Techniques for gate decomposition, routing, and mapping are continuously evolving, alongside other approaches have further explored ideas such as parameterized circuit instantiation where a feedback loop is allowed 9951320 . No standards or rules are defined to what is the fair use of these techniques for quantum devices. From the classical computing benchmarking past, we know good examples of techniques that are not allowed to encourage a fair competition of benchmark results. For instance, the definition of assertion flags is not allowed, i.e. flags that assert a certain property of the program, one that may not hold for other routines. Another example of a restriction to the configuration of a classical benchmark execution is that no more than two compilation passes without processing in between are permitted SPECHPGBenchmarksHistory . We believe on the quantum computing field there is still a need to further explore the fair usage of optimization routines and flags on transpilations to standarise the comparison of benchmark results for different devices, resembling these previous efforts on classical benchmarking. This debate extends too to the process of establishing peak and base configurations.
@@
  • Matching routines to relevant regions of the behaviour space. Even though we are still in the early ages of quantum computing, there is a huge investment to find and identify real-world applications that can benefit from this technology. This search is particularly relevant for defining which regions of the behavior space of a quantum device need to be evaluated. This effort needs to be recurrent, to prevent us from falling prey to the Goodhart’s law as warned before. We firmly believe an active effort to keep uncovering what aspects of the computational device are useful is required to keep modelling new benchmarks accordingly. These new discoveries will make some of our current benchmarks obsolete, while opening the door for future benchmarks to be created. As Hennessy and Patterson pointed out in their classic book on computing architecture, it is fallacious to assume that benchmarks retain their validity indefinitely Hennessy2011-kt .
IV.3 Standard Performance Evaluation for Quantum Computers (SPEQC)

### v4

@@ 3.4.3 Benchmarks in NISQ Era
QPack: The QPack benchmark utilizes Quantum Approximate Optimization Algorithms (QAOA) to evaluate the effectiveness of quantum computers in solving problems such as MaxCut, dominating set, and the traveling salesman problem [150]. This benchmark assesses various aspects of a quantum computer, including runtime, output accuracy, and scalability. The performance of the quantum computer is assessed by contrasting the solution obtained through QAOA execution on the quantum computer with that obtained using a classical exact algorithm. Consequently, the size of the problem that can be handle is constrained by the computational capacity of the classical computer. The QPack benchmark has been used to compare different quantum computing platforms, such as superconducting and trapped ion devices [151].
@@
Quantum LINPACK: Inspired by the classical LINPACK benchmark, the Quantum LINPACK benchmark assesses the performance of a quantum computer based on its ability to solve linear equations system using pseudo-dense random matrices. [98]. The creators of this benchmark introduce an input model known as the RAndom Circuit Block-Encoded Matrix (RACBEM), which serves as an appropriate generalization of a dense random matrix in the context of quantum computing. The success of the benchmark is measured by comparing the output vector to the ideal solution, incurring an exponential classical cost.
@@
QASMBench: The QASMBench is a benchmark that includes a set of quantum circuits and metrics for characterizing circuit properties both before and after execution on a quantum processor [96]. The set of quantum circuits are frequently used in various fields, such as chemistry, simulation, linear algebra, search algorithms, optimization, arithmetic, machine learning, fault tolerance, cryptography, and more [152]. QASMBench is divided into three categories: small-scale, medium-scale, and large-scale, each defined by the number of qubits used. As an illustration, in the small-scale scenario, the performance of quantum processes is assessed by executing quantum algorithms using 1-5 qubits such as Quantum Walk, Linear Equation, Search, and optimization, among others. In the medium-scale and large-scale scenarios, the performance of quantum processes is evaluated by running quantum algorithms using 6-15 and more than 15 qubits, respectively. The authors propose a set of circuit evaluation metrics: Circuit Width, Circuit Depth, Gate Density, Retention lifespan, Measurement Density, and Entanglement Variance, which are indicators that allow us to quantify the performance of a certain quantum circuit executed on a quantum processor. QASMBench has been assessed on IBM-Q machines, and the effectiveness of the quantum circuits is determined by computing the Hellinger distance.
@@
SupermarQ Suite: SupermarQ Suite is a benchmark that includes a set of applications that assess the performance of a quantum computer based on its ability to generate entanglement between qubits through the GHZ algorithm, demonstrate the quantum nature of the system through the Mermin-Bell test, solve combinatorial optimization problems through the quantum approximate optimization algorithm, find the ground state energy of the one-Dimensional Transverse Field Ising Model (1D TFIM) using the Variational Quantum Eigensolver algorithm (VQE) and simulate the time evolution of the 1D TFIM Hamiltonian using a quantum circuit via Trotterization for a specific number of time steps [153]. The performance of each application is assessed through different scoring function, comparing experimental outcomes with predefined ideal results. As these ideal results are known in advance, the scalability of this benchmark is ensured. The SupermarQ Suite has been employed to assess the performance of both superconducting and trapped ion devices.
@@
QED-C Benchmark: The Quantum Economic Development Consortium (QED-C) introduced the QED-C benchmark to assess the efficiency of quantum computers [114]. This benchmark comprises various algorithms, including Deutsch-Josza, Grover’s Search, VQE, Phase Estimation, Bernstein-Vazirani, Quantum Fourier Transform, and other algorithms. Unlike the Quantum Volume metric, which only uses square circuits, the QED-C benchmark incorporates circuits that vary in both depth and width. It includes circuits with high depth and low width, as well as circuits with low depth and high width. The QED-C benchmark evaluates computer performance by assessing fidelity, which involves comparing the ideal probability distribution with the experimental probability distribution. Therefore, this approach is limited by its requirement for exponentially costly classical computations to evaluate the performance of quantum computers.
@@
Chemistry simulation: The Chemistry Simulation benchmark is designed to assess the performance of quantum computers by the electronic structure calculation of molecules using variational methods [154]. The variational quantum eigensolver (VQE) algorithm has been used to calculate electronic structures using various quantum computing architectures both photonic [155], superconducting circuits [156] and trapped ions [157]. This benchmark assesses the performance of quantum computers by using ground state energy as a metric, comparing results obtained on classical computers with those derived using the VQE algorithm on quantum computers.
We have introduced the main proposals of metrics and benchmarks in the literature. Table 1 analyzes metrics, evaluating which quality attributes they fulfill according to the framework introduced in Sec. 2. We emphasize that we evaluate the quality attributes fulfilled by metrics in the NISQ era; Therefore, these evaluations are not valid for the post-NISQ era. We hope that as quantum processors evolve, benchmarks will adapt to each era.Table 1 provides an assessment of quantum computing metrics that have been proposed to measure performance-related properties of a quantum computer. A check mark indicates that the metric satisfies a given quality attribute. A cross indicates that the attribute is not satisfied in general; importantly, cross marks should be interpreted with nuance. In many cases, this may arise from limitations imposed by platform heterogeneity, strong dependence on experimental conditions, unfavorable scalability, an insufficient ability of the metric to reliably discriminate device performance, or ambiguities in the metric definition itself, among other factors. A dash mark indicates that the corresponding attribute cannot be assessed unambiguously for the metric as commonly defined, either because it is not applicable or because its evaluation depends on additional assumptions.
Figure 2: Overview of the terminology used in this work. We distinguish between techniques, quantum benchmarks, and quantum performance metrics. Representative examples of each category are provided.
Table 1: Summary of quantum computing metrics in NISQ era. In this table, we show some of the main metrics that have been proposed to measure properties related to the performance of a quantum computer, as presented in subsection 3.4.2. We assess whether these metrics satisfy the quality attribute proposed in the literature for an ideal performance metric, as discussed in subsection 2.3, although some of them are borderline. The check mark indicates that the metric satisfies the attributes; otherwise, we use the cross mark. A cross indicates that the corresponding attribute is not strictly satisfied. This may arise from limitations imposed by platform heterogeneity, strong dependence on experimental context, unfavorable scalability, an insufficient ability of the metric to reliably discriminate device performance, or ambiguities in the metric definition itself, among other factors. The dash mark indicates that it is complicated to define if the metric satisfies the criterion because we cannot characterize it formally. Performance definitions: Scalability refers to the size of the processor; Quality evaluates how accurately the quantum processor performs the tasks it was designed or programmed to execute; Speed concerns the time required to process a certain task.
@@
Number of Qubits Scalability
@@
Readout Fidelity $\times$ $\times$ Quality
(Average) Decoherence Times $\times$ $\times$ Quality
@@
Quantum Volume $\times$ $\times$ $\times$ Scalability / Quality
Q-Score $\times$ $\times$ Scalability / Quality
CLOPS $\times$ $\times$ $\times$ Speed
@@
Cross Entropy Benchmarking $\times$ - Quality
@@
4 Roadmap for Quantum Benchmarking
@@
4.1 Guidelines for the benchmark practitioner
Here, we present a set of high-level guidelines to assist benchmark designers in making informed decisions when creating benchmarks that accurately reflect the performance of the assessed quantum devices. These guidelines build upon existing industry standards and take into account insights from classical benchmarking sources (as discussed in Sec. 2), being system-independent:
@@ 4.1 Guidelines for the benchmark practitioner
Guideline 2. Benchmarks must uphold the key quality attributes: relevance, reproducibility, fairness, verifiability, and usability. In Subsec. 2.1, we introduced a set of quality attributes initially suggested as guidelines for classical benchmark designers. We encourage that quantum benchmarks be designed to satisfy these quality attributes. One of our key proposals and recommendations is that any newly defined benchmark should be accompanied by an analysis that clearly identifies which properties it satisfies and which it does not, along with a justification for the relevance or irrelevance of each property in the given context. This practice will help the community understand the appropriate contexts in which these benchmarks or metrics should be applied, as well as highlight key limitations that future developments should aim to address. As an illustrative example, we briefly analyze one of the benchmarks previously discussed in Subsec. 3.4.3: the QED-C Benchmark. This benchmark, or benchmark suite, has been continuously updated to incorporate a variety of application-oriented benchmarks for quantum computers. The final output of the benchmark can be interpreted either as a vector comprising the results of multiple specified metrics, or as a collection of individual benchmarks, each with its own set of metrics. Regardless of the interpretation, the resulting analysis remains largely consistent. In this analysis, we treat it as a benchmark suite. Some components of the suite, such as those based on Shor or Grover algorithms, are not relevant in the current technological era, a point acknowledged in the original guidelines. Others, such as benchmarks based on the Variational Quantum Eigensolver (VQE), remain pertinent. In principle, the benchmarks are reproducible, provided that the corresponding metrics are reported alongside detailed information about all parameters and device calibration. They are also fair, as they do not introduce biases tied to specific hardware architectures or artificial constraints. The methodology is well-defined and established, contributing to the overall verifiability of the suite. However, some benchmarks, such as those involving Hamiltonian simulation, do not scale efficiently to larger qubit counts due to their reliance on classical simulation. This reliance renders them unverifiable for systems with many qubits, as classical resources become insufficient to reproduce or validate the results. Moreover, not all benchmarks in the suite are usable at scale; for example, in cases like Hamiltonian simulation, the required number of measurement shots increases significantly with the number of qubits, making them unsuitable for systems with many qubits.
Guideline 3. Benchmarks must offer peak and base performance results. The use of base and peak performance is motivated by the need to enable fair comparison between devices. In addition, fixing the set of allowed optimizations promotes the reproducibility of benchmark results. Base performance is intended to capture a baseline configuration that reflects how a device performs under a common and restricted set of rules, thereby enabling meaningful comparison across architectures. This baseline corresponds to a standard, documented, and restricted configuration that a non-expert user can reasonably reproduce. In contrast, peak performance aims to capture the best achievable performance of a given platform when optimization techniques are employed, allowing innovation without constraining technological progress. Reporting both base and peak results ensures that benchmarks remain comparable while still reflecting the full potential of current quantum processors.
@@
The classical baseline debate with arguments against and for an extended use of optimization flags can be extended to quantum devices. We encourage benchmarks practitioners to follow a similar approach as the final compromise established for SPEC in January 1994: “Every SPEC CPU benchmark measurement has to measure performance with a restricted baseline set of flags, and optionally with a more extended peak set of flags”[158]. Having both metrics, peak and base, allows different vendors to share a common baseline for comparison that represents also the average non expert consumer, while enabling also the publication of the peak performance with optimised results leveraging expert knowledge over the architectures.
@@
Guideline 4. Benchmarks metrics must uphold the key quality attributes: practical, repeatable, reliable, and consistent. They should also preferably be independent and linear. As discussed in Sec. 2, for a metric to be useful in build a good benchmark it must present attributes that correlate later with quality benchmark attributes. Additionally, while linearity and independence are desirable metric attributes, they are not strictly required for effectively capturing the performance of a quantum device. Regarding benchmarks, we encourage that each proposal of a new metric includes an analysis of the properties it satisfies as well as those it does not.
@@
Figure 3: Summary of the proposed roadmap for quantum benchmarking. The figure highlights five guidelines addressing (1) the need for era-dependent benchmarking goals, (2) key benchmark quality attributes, (3) the reporting of base and peak performance, (4) desirable properties of benchmark metrics, and (5) the importance of using complementary benchmark suites rather than relying on a single test.
4.2 Benchmarking structure
@@
Device verification is a crucial process to ensure that the device is working properly. Just as vendors in classical computing certify the validity of their products before consumer release, similar steps must be taken in the quantum computing field. However, the techniques for verification differ due to the fundamentally different nature of quantum computers, rooted in quantum mechanics. Since we are in NISQ era, the concepts of verification and benchmarking overlap because we can not ensure that a quantum computer will work properly. An example of device verification method is the cross-entropy benchmarking (XEB), which has been employed to verify quantum processors based on superconducting circuits [159, 60].
@@
  • Device and routine execution verification. Post-NISQ era, a significant challenge we must address, which is becoming increasingly apparent with advancements in available quantum computers even within NISQ, is verifying the correctness of the results they produce. As mentioned earlier in Sect. 3, we know that the problem to verify if every language in BQP admits an interactive proof where the prover is in BQP too, and the verifier is in BPP is still opened [101]. This question has been thoroughly examined within the field, even extending to inquiries about the falsifiability of quantum mechanics as a theory [160]. Still, it is not the only approach explored for verifiability so far in the field, with other protocols explores with single-provers and multi-provers either quantum or classical [100].
    In practice, verification in quantum computing depends both on the computational era—such as the NISQ or post-NISQ era—and on the specific aspect of the quantum computer that is being verified. In the NISQ era, verification remains feasible for small system sizes through cross-validation with classical simulations, which compare the results generated by quantum devices with exact solutions for a small number of qubits or with approximate classical methods applicable only in restricted regimes. In addition, statistical tests can be employed to assess whether experimentally estimated quantities exceed predefined thresholds with a given level of confidence, thereby providing probabilistic guarantees even when exact verification is not possible. As quantum processors scale beyond the NISQ regime, verification becomes substantially more challenging. In the post-NISQ era, building confidence in large-scale quantum computations relies on complementary forms of indirect evidence. These include ensemble-based analyses across related problem instances, internal consistency checks based on properties expected to be preserved by the computation, robustness tests under controlled perturbations, validation against analytical results where available, and cross-verification across different quantum devices [93, 100].
    Finally, It is also important to distinguish between different notions of verification. On the one hand, verification may refer to assessing whether a given quantum circuit has been executed correctly on a device, which is typically addressed through measures related to process fidelity, implementation accuracy, or benchmarking techniques. On the other hand, verification may refer to establishing that a computation achieves a genuine quantum advantage over classical methods, which constitutes a substantially harder and more fundamental problem rooted in computational complexity. While the former is an essential component of benchmarking and device characterization, the latter generally lies beyond the scope of experimental verification. Acknowledging this distinction is crucial for the correct interpretation of benchmarking results.
@@
  • Definition of good metrics. In this study, we have introduced in Subsec. 2.3 the quality attributes that good metrics should ideally meet. Additionally, we have examined various proposals put forth by researchers in the field in Subsec. 3.4. However, we contend that further work is necessary to precisely define the relevance of the metrics. In other words, we need to determine which aspects of the devices behavior space are being captured by these measurements. Establishing a correlation between metrics and their actual relevance to performance is crucial for obtaining diverse perspectives of the device, and a thorough assessment. Throughout the NISQ era, prioritizing metrics focused on scalability, accuracy, and quality can better guide the development of robust and practically useful quantum devices. While gate speed remains a relevant and informative metric—especially given the variability in gate durations across different technological platforms—we consider it complementary to those assessing overall system performance and reliability. In our view, scalability, accuracy, and quality represent more immediate challenges for current quantum computing architectures. Therefore, speed-related metrics are valuable, but should be regarded as secondary to those that capture the system’s computational potential and robustness. However, speed may become a necessary performance indicator as we transition towards the Fault Tolerant era or implement additional partial quantum error correction methods.
@@
  • Standarise transpilation and compilation process. While we have discussed in detail the problematic regarding the transpilation, which can be mapped to the idea of how to optimise the intermediate representation (IR) language descriptor for quantum programs (e.g. OpenQASM [161] or QASM [162]), we have not detailed the mapping into the physical layer. This aspect of the compilation and transpilation funnel relates to the Quantum Physical Operations Language (QPOL) [162]. Due to the lack of a standard architecture for the development of the quantum technologies and the complex nature of simplifying the intermediate representation (IR) of circuits, much of the research focus has revolved around this aspect. In the future, we might have to include the exploration of QPOL optimization techniques through standard compilers within the industry. Even today, transpilation alone is witnessing a surge of initiatives aiming to delve deeper into its complexities. Techniques for gate decomposition, routing, and mapping are continuously evolving, alongside other approaches have further explored ideas such as parameterized circuit instantiation where a feedback loop is allowed [163]. No standards or rules are defined to what is the fair use of these techniques for quantum devices. From the classical computing benchmarking past, we know good examples of techniques that are not allowed to encourage a fair competition of benchmark results. For instance, the definition of assertion flags is not allowed, i.e. flags that assert a certain property of the program, one that may not hold for other routines. Another example of a restriction to the configuration of a classical benchmark execution is that no more than two compilation passes without processing in between are permitted [158]. We believe on the quantum computing field there is still a need to further explore the fair usage of optimization routines and flags on transpilations to standarise the comparison of benchmark results for different devices, resembling these previous efforts on classical benchmarking. This debate extends too to the process of establishing peak and base configurations.
@@
  • Matching routines to relevant regions of the behaviour space. Even though we are still in the early ages of quantum computing, there is a huge investment to find and identify real-world applications that can benefit from this technology. This search is particularly relevant for defining which regions of the behavior space of a quantum device need to be evaluated. This effort needs to be recurrent, to prevent us from falling prey to the Goodhart’s law as warned before. We firmly believe an active effort to keep uncovering what aspects of the computational device are useful is required to keep modelling new benchmarks accordingly. These new discoveries will make some of our current benchmarks obsolete, while opening the door for future benchmarks to be created. As Hennessy and Patterson pointed out in their classic book on computing architecture, it is fallacious to assume that benchmarks retain their validity indefinitely [164].
Key points of the section: Roadmap for Quantum Benchmarking What benchmarking is: In quantum computing, a benchmark is a test or test suite used to compare quantum processors by measuring performance under a specified workload. What performance means: Performance is inherently multidimensional and context-dependent, involving aspects such as execution time, throughput, resource usage, accuracy, energy consumption, and scalability. Central challenge: The choice of workload is critical—benchmarks only measure performance relative to the behaviors they exercise, and unrepresentative workloads lead to misleading results. Good benchmark principles: Useful benchmarks aim to be relevant to real use cases, reproducible, fair across systems, verifiable, and usable in practice, though these properties often involve trade-offs. Benchmark diversity: Benchmarks can differ by workload strategy (fixed-work, fixed-time, or variable), by implementation level (specification- vs. kit-based), and by realism (synthetic tests to full applications). Role of metrics: Benchmarks rely on performance metrics, which should be practical, repeatable, reliable, and consistent; no single metric captures all aspects of performance. Interpretation risks: Normalization, aggregation, and ratio-based comparisons can be manipulated, reinforcing that performance cannot be reduced to a single number. Main conclusion: There is no universal benchmark; meaningful evaluation requires multiple complementary benchmarks and careful interpretation aligned with the evaluation objective.
4.3 Standard Performance Evaluation for Quantum Computers (SPEQC)